a.  A batch of50 parts contains six defects. If two parts are drawn randomly one at a time without replacement, what is the probability that both parts are defective?

b. If this experiment is repeated, with replacement, what is the probability that both parts are defective?



1
Expert's answer
2021-06-09T14:08:58-0400

a.

"P(DD)=\dfrac{6}{50}\cdot\dfrac{6-1}{50-1}=\dfrac{3}{245}"

b.

"P(DD)=\dfrac{6}{50}\cdot\dfrac{6}{50}=\dfrac{9}{625}"